# From Radiation Dose to Cellular Dynamics: A Discrete Model for Simulating Cancer Therapy

### v1

@@ Abstract
Radiation therapy is one of the most common cancer treatments, and dose optimization and targeting of radiation are crucial since both cancerous and healthy cells are affected. Different mathematical and computational approaches have been developed for this task. The most common mathematical approach, dating back to the late 1970’s, is the linear-quadratic (LQ) model for the survival probability given the radiation dose. Most simulation models consider tissue as a continuum rather than consisting of discrete cells. While reasonable for large scale models, e.g., for human organs, any cellular scale effects become, by necessity, neglected. They do, however, influence growth, morphology, and metastasis of tumors. Here, we propose a method for modeling the effect of radiation on cells based on the mechanobiological CellSim3D simulation model for growth, division, and proliferation of cells. To model the effect of a radiation beam, we incorporate a Monte Carlo procedure into CellSim3D with the LQ model by introducing a survival probability at each beam delivery. Effective removal of dead cells by phagocytosis was also implemented. Systems with two types of cells were simulated: stiff slowly proliferating healthy cells and soft rapidly proliferating cancer cells. For model verification, the results were compared to prostate cancer (PC-3 cell line) data for different doses and we found good agreement. In addition, we simulated proliferating systems and analyzed the probability density of the contact forces. We determined the state of the system with respect to the jamming transition and found very good agreement with experiments.
@@ I Introduction
Simulation | Total # of steps ( | Write frequency () | Box side | Beam | Dose [Gy] | [Gy-1] | [Gy-2] |
Prostate | 7.25 | 2.5 | 330 | Present | 3 | 0.064 | 0.0167 |
6 | |||||||
9 | |||||||
15 | |||||||
20 | |||||||
3 | 0.241 | 0.067 | |||||
6 | |||||||
9 | |||||||
15 | |||||||
20 | |||||||
3 | 0.487 | 0.055 | |||||
6 | |||||||
9 | |||||||
15 | |||||||
20 | |||||||
Force distribution | 5.5 | 2 | 47 | Absent | 0 | 1 | 1 |
Present | 0.2333 | ||||||
0.4712 | |||||||
0.7792 | |||||||
500 | Absent | 0 | |||||
Present | 0.2333 | ||||||
0.4712 | |||||||
0.7792 |
@@ II.1 CellSim3D and inclusion of radiation beam
The CellSim3D [18] model and software for cell division was used as the starting point. CellSim3D is an open source software for mechanobiological simulations of cells populations [22]. We omit the model details here as they have been described in detail in several prior publications and since the code and model are freely available [18, 23, 24]; in particular, Supplemental Material provided in Ref. [24] has a very detailed description of the forces and interactions in the model.
In brief, CellSim3D is GPU accelerated and able to routinely handle 100,000’s of cells in 3-dimensions on a regular desktop computer equipped with standard NVidia graphics cards (a wide range has been tested as reported in Ref. [18]). The detailed derivation of the model and a description of the software can be found in Ref. [18] while the mapping between physical and model parameters is provided in Ref. [25]. The parameters used in this work are listed in Table 1 of Ref. [18] and will not be rewritten here. The model includes cell-cell adhesion as well as cell-cell and cell-medium friction, cell migration, cell division, and easily changeable elastic properties. Here, we have amended the original model by adding 1) the effect of a generic therapeutic radiation beam, 2) effective phagocytosis, and 3) the capability of simulating more than one cell type (here, healthy and cancer cells were used).
The effect of the radiation beam is implemented as a random process: it acts by picking a random number drawn for each cell. This number is then compared with the death probability . If the cells dies and cannot reproduce anymore. Other properties are left unchanged. This is essentially a simple Monte Carlo step.
@@ II.2 Linear-Quadratic (LQ) model
where is the dose delivered by the beam, while () and (), called survival parameters, describe the radiosensitivity of the tissue. The larger these parameters, the more sensitive the cells are to radiation [34]. The parameter is considered to model a single-track process meaning that the cellular DNA is damaged in a single radiation event. The parameter describes a situation of a two-track process in which the DNA is damaged by two separate radiation hits. Thus, the ratio is an important characteristic of the model, providing the relative importance of single and double events.
@@ II.3 The action of the radiation beam
In reality, dead cells undergo phagocytosis [38]. We modeled this process as follows: when phagocytosis is activated, the program progressively lowers the internal pressure of the dead cells, until the threshold defined by the user is reached. The rationale is that by decreasing the volume of the dead cells they become very soft and flexible, and free up space for the live cells to proliferate. The effect of this is essentially the same as the full removal of the dead cells but without the computational complexity required by restoring the topology of the system and neighbor tables. Phagocytosis is, in general, much faster than cellular reproduction [38] and hence the shrinking rate for the dead cells is set to be ten times the growth rate of alive cells. The factor of ten is arbitrary but crucial (although its value can be changed if/when needed): while the unit of time for the simulation is determined by the cell cycle [25], it is important to keep the events that occur within a single simulation in the right order.
In the current approach, two different cell types are present simultaneously. The first cell type represents healthy and the second one cancerous cells; we note that the type of cells is not limited to two, but can be varied as necessary. The two types of cells are identical in all parameters but Young’s modulus, which describes the cell stiffness; cancer cells are typically considered to be softer than healthy cells [39]. Cells with a higher Young’s modulus are less prone to deformation and therefore grow slowly when compared to softer cells. Slower growth results in slower division, which mirrors the well-known fact that cancer cells proliferate faster than healthy ones [40]. In prior studies this approach has been successfully applied to study tumor growth in epithelial tissues [41].
@@ III Results
80.3 | |||||
60.9 | |||||
39.5 |
The CellSim3D[18] model for mechanobiological studies of growth, division and proliferation of cells has been successfully validated against experimental data for the mitotic index and packing of epithelial cells [42], and it has been used to study tissue growth under different conditions [24, 43]. The addition of new features in CellSim3D requires two different types of validations. First, we compared the packing properties of cells to those obtained from experiments [44] for elastic spheres and found very good agreement. We compared also the contact forces and their distributions, in particular with respect to the jamming transition. The distribution of forces displays an exponential tail, in agreement with experimental data from elastic shells [44]. There are only very few force measurements in cellular systems, but Trepat et al. [45] measured traction distributions using Madin-Darby canine kidney cells, finding that the peak of the distribution can be well approximated by a Gaussian, while the tail is an exponential, in excellent agreement with our model predictions. This will be discussed in detail in Section III.1.1.
@@ III.1 Contact force probability density
Eight different simulations were executed using different sizes of the simulation box, and percentage of dead cells at the final step. The spatial configurations are labeled as open in the case of an infinite box with open boundaries and packed in the case of a rigid box filled at the level of more than 80% at the final step of the simulation (confluence or close to confluence). The percentages of alive cells at the end of the simulation were, respectively, 100% (beam absent), 80%, 60% and 40%. We reserved, in each simulation, 6% of the time steps for the system to relax after radiation delivery. Up to 60,000 cells were simulated in each case.
@@
in order to to determine the stress transfer and the state of the system with respect to the point of the jamming transition. In Eqs. 2 and 3 above, and are the respective amplitudes, the position of the peak and the width of the Gaussian.
@@ III.1.1 Analysis of the fit
The parameter values from fitting to an exponential, Eq. 3, and the survival percentages are shown in Table 2, and the probability densities of the dead cells in the six simulations with the beam active are shown in Fig. 1. Table 2 shows interesting features: First, the steepness of the tails of the distributions is higher in the open systems than in the packed ones. This is expected, since the packed system is characterized by less space per particle, thus causing a more abundant population of stronger forces. The same steepness pattern is found when comparing dead and alive cells, in both spatial configurations. There are two interpretations:
@@
We also observed that the relative errors of the dead cells are higher than those of alive cells. Moreover, the relative errors of the open space simulations are higher than those in the packed ones. These facts reflect the abundance of statistics at higher forces. In fact, the shorter the sample in the tail, the higher the probability for some bins of the histogram to contain only one or two points. The latter ones flatten the tail and augment the spread. This feature does not depend on the binning of the histogram and can be explained exactly as in the previous paragraph.
@@ III.1.2 Shape of the distributions
Figure 1 shows that the peaks of the probability distributions for the packed configuration are, in general, lower than the ones of the free distribution. This is simply caused by the longer tail, and the normalization of the distributions. Since the domain is longer and the area is fixed to unity, the peak at the mode must be lower. We notice that both tails have an exponential shape, but the cells that are in open space exhibit a steeper decay. Due to the normalization of the probability density, the cells in with the open boundary conditions are more peaked around the maximum. The difference in peak height is also found when comparing the force probability distributions of alive and dead cells in the packed vs. open graphs, as shown in Fig. 1. The shapes of the curves of cells from the same spatial configuration are basically the same.
@@
-1 | 0.064 | 0.0167 | Deweese et al. [47] |
-2 | 0.241 | 0.067 | Algan et al. [48] |
-3 | 0.487 | 0.055 | Leith et al. [49] |
@@ IV Conclusions
We have proposed and analyzed a new tool for radiation oncology simulations by representing the tissue as a set of cells, rather than a continuum material, using the recently proposed CellSim3D mechanobiological cell model. We implemented the effects of radiation beam with a customizable dose, tissue sensitivity and fractionation as a stochastic Monte Carlo process. In a first set of runs, we measured the probability density of the inter-cellular forces, finding an excellent agreement with experiments in many aspects, and some minor differences due to the celluylar nature of the system. With these measurements, we were able to determine the state of the system with respect to the jamming transition, gives information on how far the cells can move inside the tumor. As a further proof of concept, we investigated the effect of radiation on prostate tumor model, finding good qualitative, but mixed quantitative agreement with experimental data. This is likely due to the varying experimental conditions present in literature. For further development, the software can be made to include spatial and temporal inhomogeneities for the dose and tissue radiosensitivity as well as time dependence for, e.g., the dose rate. In addition, a more elaborated version of the survival probability can be used in place of Eq. (1), and cell migration can be used to simulate further dynamic effects of cancers, such as metastasis. Such extensions would allow for more quantitative simulations of fractionated dose, inhomogeneous irradiation protocols, or even an inhomogeneous radiosensitivity of the tissue, and provide useful comparisons with experiments or even predictions for actual treatments.
@@ Acknowledgements.
We thank Francesco Cordoni for a critical reading of the manuscript. MK thanks the Discovery and Canada Research Chairs Programs of the Natural Sciences and Engineering Research Council of Canada (NSERC) for financial support.References
@@
- Jones [1997] H. W. Jones, Record of the first physician to see Henrietta Lacks at the Johns Hopkins Hospital: history of the beginning of the HeLa cell line, Am. J. Obstet. Gynecol. 176, s227 (1997).

### v2

@@ Abstract
Radiation therapy is one of the most common cancer treatments, and dose optimization and targeting of radiation are crucial since both cancerous and healthy cells are affected. Different mathematical and computational approaches have been developed for this task. The most common mathematical approach, dating back to the late 1970’s, is the linear-quadratic (LQ) model for the survival probability given the radiation dose. Most simulation models consider tissue as a continuum rather than consisting of discrete cells. While reasonable for large-scale models (e.g., human organs), continuum approaches necessarily neglect cellular-scale effects, which may play a role in growth, morphology, and metastasis of tumors. Here, we propose a method for modeling the effect of radiation on cells based on the mechanobiological CellSim3D simulation model for growth, division, and proliferation of cells. To model the effect of a radiation beam, we incorporate a Monte Carlo procedure into CellSim3D with the LQ model by introducing a survival probability at each beam delivery. Effective removal of dead cells by phagocytosis was also implemented. Systems with two types of cells were simulated: stiff slowly proliferating healthy cells and soft rapidly proliferating cancer cells. For model verification, the results were compared to prostate cancer (PC-3 cell line) data for different doses and we found good agreement. In addition, we simulated proliferating systems and analyzed the probability density of the contact forces. We determined the state of the system with respect to the jamming transition and found very good agreement with experiments.
@@ I Introduction
| Simulation | Total # of steps () | Write frequency () | Box side | Dose [Gy] | [Gy-1] | [Gy-2] |
| Prostate | 7.25 | 2.5 | 330 | 3 | 0.064 | 0.0167 |
| 6 | ||||||
| 9 | ||||||
| 15 | ||||||
| 20 | ||||||
| 3 | 0.241 | 0.067 | ||||
| 6 | ||||||
| 9 | ||||||
| 15 | ||||||
| 20 | ||||||
| 3 | 0.487 | 0.055 | ||||
| 6 | ||||||
| 9 | ||||||
| 15 | ||||||
| 20 | ||||||
| Force distribution | 5.5 | 2 | 47 | 0 | 1 | 1 |
| 0.2333 | ||||||
| 0.4712 | ||||||
| 0.7792 | ||||||
| 500 | 0 | |||||
| 0.2333 | ||||||
| 0.4712 | ||||||
| 0.7792 |
@@ II.1 CellSim3D and inclusion of radiation beam
The CellSim3D [18] model and software was used as the basis for our simulations. CellSim3D is an open-source, GPU-accelerated framework for mechanobiological simulations of cell populations [21]. It represents each cell as a deformable elastic shell based on a C180 fullerene geometry, where nodes on the surface interact through bonded and non-bonded forces. The total force acting on a node is given by the equation
where the terms correspond to internal elasticity, curvature preservation, osmotic pressure, short-range repulsion, adhesion between neighboring cells, and friction with other cells and the medium [18, 22]. When a cell’s volume exceeds a threshold , it divides into two daughter cells according to a prescribed division plane, modeling mitosis.
Although detailed derivations and parameter mappings to experimental systems are provided in previous work [18, 23, 24, 22], here we emphasize that CellSim3D routinely handles aggregates of cells in three dimensions on a single desktop GPU. The model includes cell-cell adhesion, intercellular and cell-medium friction, migration, growth, and division, with easily adjustable elastic and mechanical properties. This combination makes it well suited for studying the collective mechanics of growing cell populations within physically realistic constraints.
Here, we have amended the original model by adding 1) the effect of a generic therapeutic radiation beam, 2) effective phagocytosis, and 3) the capability of simulating more than one cell type (here, healthy and cancer cells were used). The effect of the radiation beam is implemented as a random process: it acts by picking a random number drawn for each cell. This number is then compared with the death probability . If the cells dies and cannot reproduce anymore. Other properties are left unchanged. This is essentially a simple Monte Carlo step.
@@ II.2 Linear-Quadratic (LQ) model
where is the dose delivered by the beam, while () and (), called survival parameters, describe the radiosensitivity of the tissue. The larger these parameters, the more sensitive the cells are to radiation [33]. The parameter is considered to model a single-track process meaning that the cellular DNA is damaged in a single radiation event. The parameter describes a situation of a two-track process in which the DNA is damaged by two separate radiation hits. Thus, the ratio is an important characteristic of the model, providing the relative importance of single and double events. Consequently, for the -th cell, , where the label in refers to potentially different values of and for different cells. If the survival parameters are the same for all cells, so is the death probability .
@@ II.3 The action of the radiation beam
Moreover, the natural length scale is set by the division volume , with corresponding to a HeLa cell of volume m3 [37]. This mapping implies that one unit of simulation length is about m. In the simulations with the smallest box (side 47 units, mm), the final number of cells was below , resulting in a dense, confined configuration. In contrast, in the larger boxes (sides 330–500 units, i.e. 3.3–5.0 mm) the aggregates formed nearly spherical clusters of about cells that did not reach the boundaries. At a typical cellular packing fraction of , such a cluster occupies a volume of mm3, corresponding to an effective diameter of about 0.45 mm. Since clinical photon and proton beams exhibit spatial dose variations only on mm to cm scales [38], well above the size of our simulated tumors, it is a very good approximation to assume a spatially uniform radiation beam (i.e. constant dose ) throughout the simulation volume.
In reality, dead cells undergo phagocytosis [39]. We modeled this process as follows: when phagocytosis is activated, the program progressively lowers the internal pressure of the dead cells, until the threshold defined by the user is reached. The rationale is that by decreasing the volume of the dead cells they become very soft and flexible, and free up space for the live cells to proliferate. The effect of this is essentially the same as the full removal of the dead cells but without the computational complexity required by restoring the topology of the system and neighbor tables. Phagocytosis is, in general, much faster than cellular reproduction [40, 41] and hence the shrinking rate for the dead cells is set to be ten times the growth rate of alive cells. This ensures that phagocytosis (removal of dead cell volume) occurs much faster than cell division, preserving the correct temporal order of events. The factor can be adjusted if needed, but maintaining a clear separation of timescales is crucial: while the unit of time for the simulation is determined by the cell cycle [23], it is important to keep the events that occur within a single simulation in the right order.
In the current approach, two different cell types are present simultaneously. The first cell type represents healthy and the second one cancerous cells; we note that the type of cells is not limited to two, but can be varied as necessary. The two types of cells are identical in all parameters but Young’s modulus, which describes the cell stiffness; cancer cells are typically considered to be softer than healthy cells [42, 43, 44]. Cells with a higher Young’s modulus are less prone to deformation and therefore grow slowly when compared to softer cells. Slower growth results in slower division, which mirrors the well-known fact that cancer cells proliferate faster than healthy ones [45, 46]. In prior studies this approach has been successfully applied to study tumor growth in epithelial tissues [47].
@@ III Results
| 80.3 | |||||
| 60.9 | |||||
| 39.5 |
The CellSim3D[18] model for mechanobiological studies of growth, division and proliferation of cells has been successfully validated against experimental data for the mitotic index and packing of epithelial cells [48], and it has been used to study tissue growth under different conditions [22, 49]. To ensure reliability, the new features we introduced in CellSim3D were validated in two complementary ways: First, we compared the packing properties of cells to those obtained from experiments [50] for elastic spheres and found very good agreement. We compared also the contact forces and their distributions, in particular with respect to the jamming transition. The distribution of forces displays an exponential tail, in agreement with experimental data from elastic shells [50]. There are only very few force measurements in cellular systems [51, 52], but Trepat et al. [53] measured traction distributions using Madin-Darby canine kidney cells, finding that the peak of the distribution can be well approximated by a Gaussian, while the tail is an exponential, in excellent agreement with our model predictions. This will be discussed in detail in Section III.1.1.
@@ III.1 Contact force probability density
Eight different simulations were executed using different sizes of the simulation box, and percentage of dead cells at the final step. The spatial configurations are labeled as open in the case of an infinite box with open boundaries and packed in the case of a rigid box filled at the level of more than 80% at the final step of the simulation (confluence or close to confluence). In the packed setup, the box side was chosen to be small enough to mimic confined growth conditions such as within dense tissue or restricted in vitro assays. On the other side, the open setup, the box was defined so that the spheroidal cell aggregate developed without contacting the walls. This corresponds to a more physiologically relevant free-growth situation and also demonstrates that the model reproduces spheroidal growth independent of boundary effects. The percentages of alive cells at the end of the simulation were, respectively, 100% (beam absent), 80%, 60% and 40%. We reserved, in each simulation, 6% of the time steps for the system to relax after radiation delivery. Up to cells were present in the final step of the packed setup and simulated in each case.
@@
in order to to determine the stress transfer and the state of the system with respect to the point of the jamming transition. In Eqs. 3 and 4 above, and are the respective amplitudes, the position of the peak and the width of the Gaussian.
@@ III.1.1 Analysis of the fit
The fitted parameters of the exponential model [Eq. 4] together with the survival percentages are reported in Table 2. The normalized force probability densities of dead cells in the six simulations with the beam active are presented in Fig. 1, while Fig. 2 compares the corresponding distributions of alive cells in the simulation without beam to those obtained with the beam active. Table 2 shows interesting features: First, the steepness of the tails of the distributions is higher in the open systems than in the packed ones. This is expected, since the packed system is characterized by less space per particle, thus causing a more abundant population of stronger forces. The same steepness pattern is found when comparing dead and alive cells, in both spatial configurations. There are two interpretations:
@@
We also observed that the relative errors of the parameters extracted from dead-cell data are higher than those extracted from alive cells. Moreover, the relative errors of the open space simulations are higher than those in the packed ones. These facts reflect the abundance of statistics at higher forces. In fact, the shorter the sample in the tail, the higher the probability for some bins of the histogram to contain only one or two points. The latter ones flatten the tail and augment the spread. This feature does not depend on the binning of the histogram and can be explained exactly as in the previous paragraph.
@@ III.1.2 Shape of the distributions
Figure 1 illustrates that the peaks of the force probability distributions in the packed configuration are systematically lower than those in the open configuration. This difference is a direct consequence of normalization: since the packed systems exhibit longer tails, the probability mass is spread over a wider domain, lowering the height of the peak at the mode .
Both configurations display exponential tails, but the open systems decay more steeply. In the packed setup the variance is reduced, since the entire voxel of cells experiences essentially the same mechanical stress, leading to a more homogeneous distribution. In contrast, in the open spheroid, the pressure felt by cells varies strongly with position – cells near the center are under far higher stress than those at the periphery – resulting in a broader spread of forces and a correspondingly sharper peak around the most probable value. This behavior is consistent with mechanical analogies to jammed granular media, where interparticle force distributions are well-known to exhibit exponential tails [55].
In addition, Fig. 2 shows the distributions from simulations without beam exposure (black and gold curves) in contrast to the spread with different beam doses. In this case, both configurations, and especially the open one, exhibit a markedly wider range of forces. The absence of dead cells with lower internal pressure, which in the irradiated systems tend to soften and partially fold their structures, removes a mechanism that would otherwise dampen the forces. As a result, forces much higher than the peak value, in some cases exceeding it by up to a factor of 25, become apparent in the no-beam simulations.
The mechanical interpretation of these results is consistent with experimental measurements of stress and force transmission in 3D cell systems. Steinwachs et al. [51] used three-dimensional traction force microscopy to map stresses generated by single cells embedded in biopolymer networks, revealing highly anisotropic and intermittent force propagation. The results features exponential-like force distributions similar to those observed in our simulations. At the tissue scale, Dolega et al. [52] quantified the compressive stresses inside multicellular spheroids and demonstrated that the pressure increases toward the core, reducing cell proliferation and viability. These experimental findings align well with our simulated packed configurations, where the central regions sustain higher mechanical stress and show a greater fraction of dead or jammed cells.
@@
| -1 | 0.064 | 0.0167 | Deweese et al. [56] |
| -2 | 0.241 | 0.067 | Algan et al. [57] |
| -3 | 0.487 | 0.055 | Leith et al. [58] |
@@ IV Conclusions
We have proposed and analyzed a new tool for radiation oncology simulations by representing the tissue as a set of cells, rather than a continuum material, using the recently proposed CellSim3D mechanobiological cell model. We implemented the effects of radiation beam with a customizable dose, tissue sensitivity and fractionation as a stochastic Monte Carlo process. In a first set of runs, we measured the probability density of the inter-cellular forces, finding an excellent agreement with experiments in many aspects, and some minor differences due to the cellular nature of the system. With these measurements, we were able to determine the state of the system with respect to the jamming transition, getting information on how far the cells can move inside the tumor. As a further proof of concept, we investigated the effect of radiation on prostate tumor model, finding good qualitative, but mixed quantitative agreement with experimental data. This is likely due to the varying experimental conditions present in literature. For further development, the software can be made to include spatial and temporal inhomogeneities for the dose and tissue radiosensitivity as well as time dependence for, e.g., the dose rate. In addition, a more elaborated version of the survival probability can be used in place of Eq. (2), and cell migration can be used to simulate further dynamic effects of cancers, such as metastasis. Such extensions would allow for more quantitative simulations of fractionated dose, inhomogeneous irradiation protocols, or even an inhomogeneous radiosensitivity of the tissue, and provide useful comparisons with experiments or even predictions for actual treatments.
@@ Acknowledgements.
We thank Francesco Cordoni for a critical reading of the manuscript. MK thanks the Discovery and Canada Research Chairs Programs of the Natural Sciences and Engineering Research Council of Canada (NSERC) and Foundation PS for financial support.References
@@
- Zhao et al. [2008] L. Zhao, C. D. Kroenke, J. Song, D. Piwnica-Worms, J. J. H. Ackerman, and J. J. Neil, Intracellular water-specific mr of microbead-adherent cells: the hela cell intracellular water exchange lifetime, NMR in Biomedicine 21, 159 (2008).
- Semwal [2020] M. Semwal, The physics of radiation therapy, Journal of Medical Physics 45, 134 (2020).
@@
- Gregory and Pound [2011] C. D. Gregory and J. D. Pound, Cell death in the neighbourhood: direct microenvironmental effects of apoptosis in normal and neoplastic tissues, The Journal of Pathology 223, 177 (2011).
- Elliott and Ravichandran [2016] M. R. Elliott and K. S. Ravichandran, The dynamics of apoptotic cell clearance, Developmental Cell 38, 147 (2016).
- Cross et al. [2007] S. E. Cross, Y.-S. Jin, J. Rao, and J. K. Gimzewski, Nanomechanical analysis of cells from cancer patients, Nature Nanotechnology 2, 780 (2007).
- Plodinec et al. [2012] M. Plodinec, M. Loparic, C. A. Monnier, E. C. Obermann, R. Zanetti-Dallenbach, P. Oertle, J. T. Hyotyla, U. Aebi, M. Bentires-Alj, R. Y. Lim, and C. A. Schoenenberger, The nanomechanical signature of breast cancer, Nature Nanotechnology 7, 757 (2012).
@@
- Fabrikant et al. [1969] J. I. Fabrikant, C. L. Wisseman, and M. J. Vitak, The kinetics of cellular proliferation in normal and malignant tissues, Radiology 92, 1309 (1969), pMID: 4892837, https://doi.org/10.1148/92.6.1309 .
@@
- Steinwachs et al. [2016] J. Steinwachs, C. Metzner, K. Skodzek, N. Lang, I. Thievessen, C. Mark, S. Münster, K. E. Aifantis, and B. Fabry, Three-dimensional force microscopy of cells in biopolymer networks, Nature Methods 13, 171 (2016).
- Dolega et al. [2017] M. E. Dolega, M. Delarue, F. Ingremeau, J. Prost, A. Delon, and G. Cappello, Cell-like pressure sensors reveal increase of mechanical stress towards the core of multicellular spheroids under compression, Nature Communications 8, 14056 (2017).
@@
- Majmudar and Behringer [2005] T. S. Majmudar and R. P. Behringer, Contact force measurements and stress-induced anisotropy in granular materials, Nature 435, 1079 (2005).
@@
- Baka et al. [2023] Z. Baka, C. Godier, L. Lamy, A. Mallick, V. Gribova, A. Figarol, L. Bezdetnaya, A. Chateau, Z. Magne, M. Stiefel, D. Louaguef, P. Lavalle, E. Gaffet, O. Joubert, and H. Alem, A coculture based, 3d bioprinted ovarian tumor model combining cancer cells and cancer associated fibroblasts, Macromolecular Bioscience 23, 2200434 (2023).